I am not able to resolve the following question.
"The sum of the ages of a father and his son is 52. 3 years ago, father's age is two years more than thrice the sons age. Find the age of the son 2 years hence."
Please someone explain it.

Answer by math_helper(2461)

Say
F = Father's age now
S = Son's age now
—
You have:
F + S = 52 (eq 1) ("The sum of the ages of a father and his son is 52")
And you are given:
F - 3 = 2 + 3(S-3) (eq 2) ("3 years ago, father's age is two years more than thrice the sons age")
Specifically:
The -3 on each side is because it was 3 years ago
The 3(S-3) is "thrice the son's age 3 years ago"
and the +2 on the RHS is to comply with the father's age being "two years more than…"
—
So that's two equations in two unknowns, it should be solvable.
—
I take "Find the age of the son 2 years hence" to mean "Find the son's age two years from now. "
So once S is found, add 2 to it.
—
Solving: I will take the very first step, then you can finish:
Eq 1 tells you F=52-S
Thus eq 2 can be re-written:
(52-S)-3 = 2+3(S-3)
This can be simplified and solved for S.